Circular No. 7817

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 B1 (LINEAR) An object reported as asteroidal in appearance by LINEAR, and posted

on the NEO Confirmation Page, has been found to be cometary on CCD images taken by M. Tichý and J. Tichá at Kleť (coma diameter 7′′-8′′, on Jan. 29.8 and Feb. 1.8 UT; coma diameter 9′′ with m_{1} = 17.7–18.0 and a faint tail in p.a. 165^{o} on Feb. 4.8) and by A. Galád and L. Kornoš at Modra (‘slightly diffuse’ on Feb. 1 and 2).

2002 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 26.09423  3^{h}06^{m}22.^{s}85  +64^{o}47′16′′.2  18.2

Additional astrometry (including LINEAR prediscovery observations made on 2001 Nov. 8), the following orbital elements, and an ephemeris appear on MPEC 2002-C20.

Epoch = 2002 May 6.0 TT

T = 2002 Apr. 20.15557 TT  ω = 76.^{o}25784
e = 0.7706525  Ω = 58.16584 }2000.0
q = 2.2692128 AU  i = 50.99090

a = 9.8942127 AU n^{o} = 0.03166884 P = 31.122 years

SUPERNOVA 2002ap IN M74 E. Berger, S. R. Kulkarni, and D. A. Frail, California Institute of

Technology, report that an observation of SN 2002ap with the Very Large Array at 8.46 GHz on Feb. 1.03 UT reveals a radio source with a flux density of 375 ± 30 μJy coincident with the supernova (IAUC 7816; GCN Circ. No. 1237): “The radio emission peaks at ≈ 5 GHz and appears to be fading with flux densities in the 8.46-GHz band of 255 ± 44 μJy on Feb. 1.93 and 201 ± 47 μJy on Feb. 2.79. SN 2002ap is thus > 3 orders of magnitude less luminous in the radio band than SN 1998bw.”

SUPERNOVA 2002I IN IC 4229 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002I (cf. IAUC 7799), obtained by M. Calkins on Jan. 22.53 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova several weeks past maximum.

2002 February 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green